# Production of high-spin $\omega_{J}/\rho_{J}$ ($J=3,4,5$) mesons in $\pi^{-}p$ reactions

### v1

@@ III Production of higher-spin $\omega_{J}/\rho_{J}$ ($J=4,5$) mesons in $\pi^{-}p$ reactions
$\alpha_{R}(t)=1+\alpha^{\prime}_{R}\left(t-m_{R}^{2}\right),\qquad R=\rho,\;a_{1},\;b_{1},\;h_{1},$ (29)
with slopes $\alpha^{\prime}_{\rho}=0.8\;\mathrm{GeV}^{-2}$ Guidal et al. (1997), $\alpha^{\prime}_{a_{1}}=0.8\;\mathrm{GeV}^{-2}$ (determined from the mass of the $a_{3}(1875)$ state following Ref. Galata (2011)), $\alpha^{\prime}_{b_{1}}=0.7\;\mathrm{GeV}^{-2}$ Wang et al. (2025), and $\alpha^{\prime}_{h_{1}}=0.7\;\mathrm{GeV}^{-2}$ Wang et al. (2025).
With the amplitudes (25)–(28) and the differential cross-section formula in Eq. (14), we can now compute the total and differential cross sections for the four higher-spin states.
Figure 9: The differential cross section $d\sigma/d\cos\theta$ of the $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, $\rho_{5}(2350)$ production at different c.m. energies W = 3.5, 4, 5 and 6 GeV.
Total cross sections. The total cross sections for meson–nucleon production of the $J=4$ states $\omega_{4}(2250)$ and $\rho_{4}(2230)$, and the $J=5$ states $\omega_{5}(2250)$ and $\rho_{5}(2350)$ via $t$-channel exchange are shown in Fig. 8 as functions of the center-of-mass energy $\sqrt{s}$. The solid curves correspond to a central cutoff $\Lambda_{t}=3.5~\mathrm{GeV}$, with the shaded bands representing the uncertainty from varying $\Lambda_{t}$ by $\pm 0.5~\mathrm{GeV}$.
@@
Differential cross sections. The angular distributions are presented in Fig. 9, which shows $d\sigma/d\cos\theta$ for the four states at center-of-mass energies $W=3.5$, $4.0$, $5.0$, and $6.0\;\mathrm{GeV}$.
All panels in Fig. 9 share two key features characteristic of $t$-channel processes: (i) a very strong forward peak ($\cos\theta\to 1$) at all energies, and (ii) the forward peak becomes narrower, and its cross section value gets smaller everywhere, as the c.m. energies W increases. This reflects the kinematics of meson exchange, where small momentum transfer favors scattering at very small angles, an effect that becomes more pronounced at higher energies.
A closer inspection reveals the impact of spin. The $J=4$ state $\omega_{4}(2250)$ [Fig. 9(a)] reaches differential cross sections up to about $10\;\mu\mathrm{b}$ in the extreme forward direction, while its isospin partner $\rho_{4}(2230)$ [Fig. 9(b)] shows a similar shape but with a slightly gentler slope. The $J=5$ states [Figs. 9(c) and (d)] exhibit the expected suppression. For $\rho_{5}(2350)$, the forward differential cross section is about $0.01\ \mu\mathrm{b}$, which is still within a potentially measurable range.
@@ IV Summary
In this work, we have performed a systematic study of the production of high-spin $\omega_{J}$ and $\rho_{J}$ mesons ($J=3,4,5$) in $\pi^{-}p$ scattering, employing an effective Lagrangian approach. Our primary goal was to provide theoretical predictions for the production dynamics of these states, which are far less explored than their decay properties.
Our analysis is based on a unified framework that considers only the dominant $t$-channel meson-exchange mechanisms. For each channel, the exchanged meson (e.g., $\rho$, $\pi$, $a_{1}$, $b_{1}$, or $h_{1}$) is selected according to the $\pi$-containing decay modes predicted by the QPC model. For the well-established $J=3$ states $\omega_{3}(1670)$ and $\rho_{3}(1690)$, the model has a single adjustable parameter: the form-factor cutoff $\Lambda_{t}$. A fit to the available total and differential cross-section data yields $\Lambda_{t}=3.5\pm 0.5\ \text{GeV}$. Using this determined parameter, we subsequently made parameter-free predictions for the production of their higher-spin partners $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, and $\rho_{5}(2350)$. The high-energy behavior is consistently described by applying Reggeization to the $t$-channel propagators.
@@
    The consistent description of six different resonance production channels with a single universal parameter demonstrates the robustness of the model. It provides a coherent picture linking the decay properties (from the QPC model) to the production dynamics in scattering experiments.
In conclusion, this work provides the first comprehensive set of theoretical predictions for the production cross sections of the higher-spin $\omega_{4}$, $\rho_{4}$, $\omega_{5}$, and $\rho_{5}$ states in $\pi^{-}p$ reactions. The results offer concrete, testable guidance for future experimental searches at facilities such as J-PARC and COMPASS, particularly in the forward-angle region where the cross sections are largest. This analysis bridges spectroscopy and production studies, supporting future meson-beam experiments in completing the spectrum of light hadrons.
@@ Acknowledgements.
The explicit expressions for the projection operators of fields from spin3-3 to spin5-5 are given as follows Döring et al. (2026).

### v2

@@ III Production of higher-spin $\omega_{J}/\rho_{J}$ ($J=4,5$) mesons in $\pi^{-}p$ reactions
$\alpha_{R}(t)=1+\alpha^{\prime}_{R}\left(t-m_{R}^{2}\right),\qquad R=\rho,\;\omega,\;a_{1},\;b_{1},\;h_{1},$ (33)
with slopes $\alpha^{\prime}_{\rho}=0.8\;\mathrm{GeV}^{-2}$ Guidal et al. (1997), $\alpha^{\prime}_{\omega}=0.9\;\mathrm{GeV}^{-2}$ Guidal et al. (1997), $\alpha^{\prime}_{a_{1}}=0.8\;\mathrm{GeV}^{-2}$ (determined from the mass of the $a_{3}(1875)$ state following Ref. Galata (2011)), $\alpha^{\prime}_{b_{1}}=0.7\;\mathrm{GeV}^{-2}$ Wang et al. (2025), and $\alpha^{\prime}_{h_{1}}=0.7\;\mathrm{GeV}^{-2}$ Wang et al. (2025).
With the amplitudes (Eqs. (29)–(32)) and the differential cross-section formula in Eq. (14), we can now compute the total and differential cross sections for the six states.
Figure 9: The differential cross section $d\sigma/d\cos\theta$ of the $\omega_{2}(1975)$, $\rho_{2}(1940)$, $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, $\rho_{5}(2350)$ production at different c.m. energies W = 3.5, 4, 5 and 6 GeV.
Total cross sections. The total cross sections for meson–nucleon production of the $J=2$ states $\omega_{2}(1975)$ and $\rho_{2}(1940)$, the $J=4$ states $\omega_{4}(2250)$ and $\rho_{4}(2230)$, and the $J=5$ states $\omega_{5}(2250)$ and $\rho_{5}(2350)$ via $t$-channel exchange are shown in Fig. 8 as functions of the center-of-mass energy $\sqrt{s}$. The solid curves correspond to a central cutoff $\Lambda_{t}=3.5~\mathrm{GeV}$, with the shaded bands representing the uncertainty from varying $\Lambda_{t}$ by $\pm 0.5~\mathrm{GeV}$.
@@
Differential cross sections. The angular distributions are presented in Fig. 9, which shows $d\sigma/d\cos\theta$ for the six states at center-of-mass energies $W=3.5$, $4.0$, $5.0$, and $6.0\,\mathrm{GeV}$.
All panels in Fig. 9 share two key features of $t$-channel processes: (i) a very strong forward peak ($\cos\theta\to 1$) at all energies, and (ii) the forward peak narrows with increasing energy. This reflects the kinematics of meson exchange, where small momentum transfer favors scattering at very small angles.
A closer inspection reveals significant effects of both spin and isospin. For the $J=2$ resonances, a clear isospin splitting is observed. The $\omega_{2}(1975)$ [Fig. 9(a)] exhibits forward differential cross sections on the order of $10\,\mu\mathrm{b}$ across the measured energy range. In contrast, its isospin partner, the $\rho_{2}(1940)$ [Fig. 9(d)], displays a consistently larger magnitude, with forward cross sections on the order of $10^{2}\,\mu\mathrm{b}$. This indicates that the production of the isovector $\rho_{2}$ state is significantly enhanced compared to the isoscalar $\omega_{2}$ state for $J=2$.
The $J=4$ state $\omega_{4}(2250)$ [Fig. 9(b)] reaches differential cross sections of up to about $10;\mu\mathrm{b}$ in the extreme forward direction, while its isospin partner $\rho_{4}(2230)$ [Fig. 9(e)] shows a similar shape but with a slightly gentler slope and a lower peak magnitude. The $J=5$ states [Figs. 9(c) and (f)] exhibit the expected strong suppression, particularly for the $\rho_{5}(2350)$ state.
@@ IV Summary
In this work, we have performed a systematic study of the production of high-spin $\omega_{J}$ and $\rho_{J}$ mesons ($J=2,3,4,5$) in $\pi^{-}p$ scattering, employing an effective Lagrangian approach. Our primary goal was to provide theoretical predictions for the production dynamics of these states, which are far less explored than their decay properties.
Our analysis is based on a unified framework that considers only the dominant $t$-channel meson-exchange mechanisms. For each channel, the exchanged meson (e.g., $\rho$, $\pi$, $\omega$, $a_{1}$, $b_{1}$, or $h_{1}$) is selected according to the $\pi$-containing decay modes predicted by the QPC model. For the well-established $J=3$ states $\omega_{3}(1670)$ and $\rho_{3}(1690)$, the model has a single adjustable parameter: the form-factor cutoff $\Lambda_{t}$. A fit to the available total and differential cross-section data yields $\Lambda_{t}=3.5\pm 0.5\ \text{GeV}$. Using this determined parameter, we subsequently made parameter-free predictions for the production of their lower-spin and higher-spin partners $\omega_{2}(1975)$, $\rho_{2}(1940)$, $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, and $\rho_{5}(2350)$. The high-energy behavior is consistently described by applying Reggeization to the $t$-channel propagators.
@@
    The consistent description of eight different resonance production channels with a single universal parameter demonstrates the robustness of the model. It provides a coherent picture linking the decay properties (from the QPC model) to the production dynamics in scattering experiments.
In conclusion, this work provides the first comprehensive set of theoretical predictions for the production cross sections of the high-spin $\omega_{2}$, $\rho_{2}$, $\omega_{4}$, $\rho_{4}$, $\omega_{5}$, and $\rho_{5}$ states in $\pi^{-}p$ reactions. The results offer concrete, testable guidance for future experimental searches at facilities such as J-PARC and COMPASS, particularly in the forward-angle region where the cross sections are largest. This analysis bridges spectroscopy and production studies, supporting future meson-beam experiments in completing the spectrum of light hadrons.
@@ Acknowledgements.
The explicit expressions for the projection operators of fields from spin3-3 to spin5-5 are given as follows Döring et al. (2026); Kim et al. (2025).
$\Delta_{\alpha_{1}\alpha_{2}}^{\beta_{1}\beta_{2}}(2,p)=\frac{1}{2}\left(\bar{g}_{\alpha_{1}}^{\beta_{1}}\bar{g}_{\alpha_{1}}^{\beta_{2}}+\bar{g}_{\alpha_{1}}^{\beta_{1}}\bar{g}_{\alpha_{2}}^{\beta_{1}}-\frac{2}{3}\bar{g}_{\alpha_{1}\alpha_{2}}\bar{g}^{\beta_{1}\beta_{2}}\right).$ (35)